# MCDFN: Supply Chain Demand Forecasting via an Explainable Multi-Channel Data Fusion Network Model Integrating CNN, LSTM, and GRU

### v3

@@ Abstract
Accurate demand forecasting is vital for optimizing supply chain management and enhancing organizational resilience. Traditional forecasting methods, relying on simple arithmetic, often fail to capture complex patterns caused by seasonal variability and special events. Although deep learning techniques have advanced, the lack of interpretable models hampers understanding and explaining predictions. We introduce the Multi-Channel Data Fusion Network (MCDFN), a novel hybrid deep learning architecture integrating multiple data modalities for superior demand forecasting. MCDFN utilizes Convolutional Neural Networks, Long Short-Term Memory networks, and Gated Recurrent Units to extract spatial and temporal features from time series data. Comparative benchmarking against seven other deep-learning models validates MCDFN’s efficacy, showing it outperforms its counterparts across key metrics with a mean squared error of 23.5738, root mean squared error of 4.8553, mean absolute error of 3.9991, and mean absolute percentage error of 20.1575%. A paired t-test with a p-value of 5% indicated no significant difference between MCDFN’s predictions and actual values. A 10-fold cross-validated statistical paired t-test further confirmed its robustness, significantly outperforming other models at a 5% significance level. To address the “black box” nature of MCDFN, we employ explainable AI techniques such as ShapTime and Permutation Feature Importance, offering insights into model decision-making processes. This research advances demand forecasting methodologies and provides practical guidelines for integrating MCDFN into existing supply chain systems while also outlining future research directions to enhance scalability, integrate external factors for better predictions, and develop user-friendly interfaces for real-world applications.
@@ 1 Introduction
Integrating one’s SC is a critical aspect of today’s industry, and it’s one that all upstream companies have come to understand, thanks to the growing relevance of SC integration. As a result of this type of integration, information is often shared between numerous business partners (Carbonneau et al.,, 2007). Information sharing and integration across the SC are increasingly important to businesses. Even though these programs reduce forecast mistakes, they are not universal or complete, and forecast errors continue to exist (Carbonneau et al.,, 2008). In supply chain management (SCM), demand is crucial information that may be shared and utilized. Demand projections that are accurate and trustworthy are essential for SC managers to use in their planning and decision-making processes. Many SC choices are based on demand forecasting, including demand planning and order fulfillment, production planning, and inventory management (Abolghasemi et al.,, 2020).
It is well acknowledged that information exchange throughout the SC is the most effective method of preventing demand signal distortions. Even though integrated SCs are enticing, implementing them is not always practical. Many factors could make such long-term, stable joint undertakings more difficult. Premkumar lists several essential challenges that must be solved to allow for successful SC collaboration. Lengthy relationship management, unwillingness to share information, the complexity of managing an extensive SC, the ability of employees to support SC, and performance evaluation and incentives to encourage SC are just a few of the factors to consider. Despite several attempts to enable successful extended SC collaboration, most firms have failed to solve these challenges, and it is unlikely they will do so soon. More importantly, in many SCs, power regimes and sub-regimes can obstruct SC optimization (Cox et al.,, 2001; Watson,, 2001), which can be especially damaging to SC performance. Furthermore, power regimes and sub-regimes can impede SC optimization in many supply networks. So, even if the integration of systems and information sharing is technically viable organizationally, it may not be feasible due to considerable disruptions in the structure of power (Premkumar,, 2000). The bullwhip effect is magnified when SC participants share advanced knowledge.
@@
Information distortion is one of the troublesome and pricey phenomena in SC. Better collaboration and accurate forecasting are required to reduce the bullwhip effect resulting from this information distortion. In this research, our objective is to forecast future demand using hybrid DL techniques. It can boost the organization’s service level and enable better scheduling and planning. Therefore, it will reduce the cost of the organization to a beneficial extent.
Despite significant advancements in the application of DL and ML for demand forecasting in SCM, several research gaps persist that we aim to address in this paper. Specifically, there is a pressing need for more robust and scalable DL models tailored to handle the complexities and dynamic nature of SCs. Existing models often encounter challenges related to scalability and robustness, particularly when applied to large, real-world datasets. Although the current research trend in AI has witnessed significant improvements with the emergence of generative AI, the utilization of hybrid and recent interpretable DL models remains largely unexplored in the domain of SC demand forecasting. Moreover, current DL models often lack full integration of external variables such as economic indicators, weather conditions, and special events, which are known to have a significant impact on demand. Incorporating domain-specific knowledge and external factors can enhance the predictive power of these models. Another significant challenge with DL models is their inherent “black box” nature, which can impede trust and adoption among practitioners. Therefore, enhancing the interpretability and transparency of these models is crucial for facilitating their practical utility. Although hybrid models that combine different DL techniques have shown promise, there is still ample room for improvement in their application to multi-horizon forecasting tasks. By addressing these gaps, this research endeavors to advance the state of DL in demand forecasting, ultimately providing more accurate, interpretable, and actionable insights for SCM practitioners.
@@
    We introduce a novel neural network architecture called the Multi-Channel Data Fusion Network (MCDFN), designed to analyze time series data comprehensively by incorporating multiple parallel channels within the network.
@@
    This study shows no significant difference between the actual and predicted values of MCDFN through a paired t-test for each model using a p-value of 5%.
@@ 2 Literature Review
The application of DL in demand forecasting has seen significant advancements, especially in SCM. These innovations aim to enhance predictive accuracy and operational efficiency, which is crucial for maintaining robust and resilient SCs. This literature review explores recent studies from 2017 to 2024, highlighting the evolving methodologies and their implications in SCM. Recent research has introduced Temporal Fusion Transformers (TFT) as a powerful tool for handling multi-horizon forecasting problems (Lim et al.,, 2021). TFT integrates gating mechanisms, variable selection networks, and attention-based models to enhance accuracy and interpretability. This model has demonstrated superior performance across various applications by effectively managing static and dynamic inputs and providing insights into feature importance and relevant time steps. Terrada et al. discussed the integration of AI within SCM 4.0, emphasizing the role of DL methods such as Auto-Regressive Integrated Moving Average (ARIMA) and LSTM (Terrada et al.,, 2022). Their findings show that DL methods significantly improve demand forecasting accuracy using historical transaction data, thus enhancing decision-making processes and balancing supply and demand more effectively.
Attention-based models have shown great promise in demand forecasting by capturing long-term dependencies and relevant features from time series data. Google Research demonstrated that attention mechanisms could improve traditional DL models’ performance by focusing on the most critical parts of the input data, thereby increasing forecast accuracy and interoperability (Arık and Pfister,, 2021). Techniques such as LIME (Local Interpretable Model-agnostic Explanations) and SHAP (SHapley Additive exPlanations) have been applied to address the need for transparency in DL models. These methods help explain the predictions by showing the contribution of individual features, making DL models more user-friendly and trustworthy.
@@ 3.1 Data Description and Preprocessing
    Timestamp column: This column records the exact date of each sale, providing a continuous chronological sequence vital for time series analysis. The timestamps are formatted in a standard date format (YYYY-MM-DD), facilitating easy parsing and manipulation for analytical purposes.
@@
    Quantity sold column: This column reflects the units sold on each corresponding day. The sales quantities are recorded as integers, representing the daily demand for the product.
Key observations and preprocessing steps are discussed below.
@@
3.1.1 Time Effects
@@
Product demand exhibits seasonality, being lower at the beginning and end of each year and peaking in the middle. To represent this, the following DateTime features are extracted: ‘Day of Month,’ ‘Week,’ ‘Month,’ and ‘Year’. We incorporated the conversion of the DateTime feature into cyclical features to handle the periodic nature of time series data more effectively. This conversion is crucial for capturing the seasonal patterns inherent in the data, which traditional linear models might fail to identify adequately. The transformation process generates sine and cosine signals for various time-based components, such as timestamp, month, day, week, and year. This approach ensures that the periodicity is represented in a form that machine learning models can interpret more accurately. ‘Day of Month’ and ‘Month’ features are transformed into cyclic features using equations 1, 2, 3, and 4.
@@
3.1.2 Effect of holidays
Some products have higher sales during holidays. An ‘is_holiday’ feature was included to capture periods of potentially higher sales.
3.1.3 Standardization
Ensuring proper feature scaling is imperative in the preparatory stages of NN training. In sales forecasting, there is a scale difference between sales and other features. A prevalent method employed for this purpose is normalization, wherein the mean and standard deviation of each feature except for the ‘is_holiday’ feature are subtracted and divided, respectively. This procedure aids in standardizing the range of feature values, thereby enhancing the efficiency of network convergence. The ‘is_holiday’ feature is already in binary format, requiring no further normalization. Consequently, accurate training cannot be performed when inputs are used without being normalized. Moreover, sales follow Gaussian distribution inherently. After reviewing available normalization methods, the sale feature is standardized using the following equation.
@@ 3.2 Deep Learning Models
In multi-step prediction tasks, the modeling paradigm shifts from forecasting individual future points to predicting a sequence of future values, requiring a more sophisticated approach. Unlike single-step forecasting, which predicts only a single future point, a multi-step model generates a coherent sequence of future observations. A prominent strategy for this is the “single-shot” model, where the entire sequence prediction is formulated and executed in a single inference step. In our research, we adopted the single-shot prediction methodology, enabling the model to forecast the complete time series trajectory simultaneously. This approach ensures that the demands across all output time steps are predicted concurrently, fostering a comprehensive understanding of the temporal dynamics in the dataset. Specifically, our multi-step models are trained using daily samples as input data to predict 30 days into the future based on a preceding window of 30 days’ historical data (Figures 6 and 7). We developed and benchmarked MCDFN against seven fine-tuned state-of-the-art DL models in SC demand forecasting, including BiLSTM, CNN, RNN, Vanilla LSTM, Stacked LSTM, MLP, and GRU.
@@ 3.2.1 BiLSTM
The activation functions, namely the hyperbolic tangent (tanh) and the sigmoid function for recurrent connections, imbue the model with non-linearity and memory retention capabilities crucial for sequence modeling. A dropout layer with a dropout rate of 0.2 is employed after the LSTM layer to mitigate overfitting and enhance generalization performance. Following sequence processing, a flattening operation transforms the output into a one-dimensional vector, facilitating subsequent dense layer operations. A densely connected layer with an output dimensionality of 30 is introduced, initializing the kernel weights with zeros. This layer maps the extracted temporal features into the desired output space, effectively reconstructing the predicted sequence.
@@ 3.2.5 Stacked LSTM
The stacked LSTM model was designed using TensorFlow’s Keras API, which begins with an input layer for data shaped $(30,10)$, followed by two LSTM layers. Each LSTM layer is configured with 512 units, which was determined through hyperparameter tuning and employs ‘tanh’ activation and ‘sigmoid’ recurrent activation functions. Both LSTM layers use the ‘return_sequences=True’ parameter to pass the sequence output to the next layer. Between the LSTM layers, a Dropout layer is included to mitigate overfitting, with 0 dropout rate after tuning. After the LSTM layers, the model output is flattened and passed through a Dense layer, initialized with zeros, ensuring unbiased predictions. Finally, the output is reshaped to the desired dimensions (30, 1). Stacked LSTM was evaluated and achieved the best validation mean squared error of 28.17 after 58 trials, demonstrating its ability to capture complex temporal dependencies and provide accurate demand forecasts for an output sequence length of 30 steps.
3.2.6 MLP
Our research utilized a Multilayer Perceptron (MLP) model for demand forecasting, implemented using TensorFlow’s Keras API. The model starts with an input layer to handle data shaped $(30,10)$, followed by a dense layer with 512 units chosen through hyperparameter tuning. The activation function for this layer was selected as ‘tanh’ from the options ‘relu’ and ‘tanh’. A Dropout layer with a rate of 0.0 is included to prevent overfitting. Finally, the model outputs predictions through a dense layer with a linear activation function to match the dimensionality of the target feature. The MLP model was evaluated and achieved the best validation mean squared error of 27.79 after 54 trials.
The input layer of an MLP receives the input vector $\mathbf{x}$. For each hidden layer $l$ (where $l=1,2,\ldots,L$ with $L$ being the number of hidden layers), the output $\mathbf{h}^{(l)}$ is computed using equation 15.

### v5

@@ Abstract
Accurate demand forecasting is vital for optimizing supply chain management and enhancing organizational resilience. Traditional forecasting methods, relying on simple arithmetic, often fail to capture complex patterns caused by seasonal variability and special events. Although deep learning techniques have advanced, the lack of interpretable models hampers understanding and explaining predictions. We introduce the Multi-Channel Data Fusion Network (MCDFN), a novel hybrid deep learning architecture integrating multiple data modalities for superior demand forecasting. MCDFN utilizes Convolutional Neural Networks (CNNs), Long Short-Term Memory networks (LSTMs), and Gated Recurrent Units (GRUs) to extract spatial and temporal features from time series data. Comparative benchmarking against seven other deep-learning models validates MCDFN’s efficacy, showing it outperforms its counterparts across key metrics with a mean squared error (MSE) of 23.5738, root mean squared error (RMSE) of 4.8553, mean absolute error (MAE) of 3.9991, and mean absolute percentage error (MAPE) of 20.1575%. Theil’s U statistic of 0.1181 ($U<1$) of MCDFN indicates its superiority over the naive forecasting approach, and a 10-fold cross-validated statistical paired t-test with a p-value of 5% indicated no significant difference between MCDFN’s predictions and actual values. To address the “black box” nature of MCDFN, we employ explainable AI techniques such as ShapTime and Permutation Feature Importance, offering insights into model decision-making processes. This research advances demand forecasting methodologies and provides practical guidelines for integrating MCDFN into existing supply chain systems.
@@ 1 Introduction
Integrating one’s SC is a critical aspect of today’s industry, and it’s one that all upstream companies have come to understand, thanks to the growing relevance of SC integration. As a result of this type of integration, information is often shared between numerous business partners (Carbonneau et al.,, 2007). Information sharing and integration across the SC are increasingly important to businesses. Even though these programs reduce forecast mistakes, they are not universal or complete, and forecast errors continue to exist (Carbonneau et al.,, 2008). In SC management (SCM), demand is crucial information that may be shared and utilized. Demand projections that are accurate and trustworthy are essential for SC managers to use in their planning and decision-making processes. Many SC choices are based on demand forecasting, including demand planning and order fulfillment, production planning, and inventory management (Abolghasemi et al.,, 2020).
It is well acknowledged that information exchange throughout the SC is the most effective method of preventing demand signal distortions. Even though integrated SCs are enticing, implementing them is not always practical. Many factors could make such long-term, stable joint undertakings more difficult. Premkumar lists several essential challenges that must be solved to allow for successful SC collaboration (Premkumar,, 2001). Lengthy relationship management, unwillingness to share information, the complexity of managing an extensive SC, the ability of employees to support SC, and performance evaluation and incentives to encourage SC are just a few of the factors to consider. Despite several attempts to enable successful extended SC collaboration, most firms have failed to solve these challenges, and it is unlikely they will do so soon. More importantly, in many SCs, power regimes and sub-regimes can obstruct SC optimization (Cox et al.,, 2001; Watson,, 2001), which can be especially damaging to SC performance. Furthermore, power regimes and sub-regimes can impede SC optimization in many supply networks. So, even if integrating systems and information sharing is technically viable organizationally, it may not be feasible due to considerable disruptions in the power structure (Premkumar,, 2000). The bullwhip effect is magnified when SC participants share advanced knowledge.
@@
Information distortion is one of the troublesome and pricey phenomena in SC. Better collaboration and accurate forecasting are required to reduce the bullwhip effect resulting from this information distortion. In this research, our objective is to forecast future demand using hybrid DL techniques. It can boost the organization’s service level and enable better scheduling and planning. Therefore, it will reduce the organization’s costs to a beneficial extent. Improved demand forecasting offers considerable economic advantages by enhancing cost efficiency across various operational dimensions. Studies reveal that refined forecasting practices can yield operational cost savings of around 20%, as more precise demand estimates allow firms to streamline their operations and eliminate unnecessary expenditures. In the area of inventory management, more accurate forecasts enable companies to achieve savings between 5% and 10% on transportation, storage, and production planning while also reducing inventory holding costs and minimizing the risk of stockouts (Kogler et al.,, 2021; Sagaert et al.,, 2018). Additionally, by aligning inventory more closely with actual demand, improved forecasting can decrease backorder rates by up to 30%, thereby enhancing customer satisfaction and lowering costs associated with expedited shipping and lost sales (Salehzadeh et al.,, 2020). Moreover, avoiding forecasting inaccuracies is essential to mitigate opportunity costs, as even a modest 10% stockout rate can result in significant revenue losses, underscoring the pivotal role of economic precision in demand forecasting for maintaining competitive market performance (Badr and Ahmed,, 2023; Mohsen,, 2023).
Despite significant advancements in applying DL and ML to demand forecasting in SCM, several key challenges remain. One major gap is the need for more robust and scalable DL models that can effectively handle the complexities and dynamic nature of SCs. Existing models often struggle with scalability and robustness, especially when applied to large, real-world datasets. Traditional demand forecasting methods, such as autoregressive integrated moving averages (ARIMA) and exponential smoothing, face difficulties in capturing non-linearity and complex seasonality in SCs. While ML models like random forests and gradient boosting improve forecasting accuracy, they fail to handle heterogeneous data sources effectively. DL models, including LSTMs and transformers, offer further improvements but often function as black boxes, limiting interpretability and trust among practitioners. Enhancing model transparency is crucial for practical adoption in SCM. Although hybrid DL models have shown promise, there is still room for improvement, particularly in multi-horizon forecasting. To address these gaps, we propose Multi-Channel Data Fusion Network (MCDFN) as a hybrid approach integrating multi-channel parallel data inputs. It leverages CNNs for spatial feature extraction, LSTMs for capturing long-term dependencies, and GRUs for efficient sequence modeling. Additionally, its model-agnostic explainability module mitigates the “black-box” issue by providing interpretable feature importance metrics, making it a more practical solution for SC decision-making.
@@
    We introduce a novel hybrid neural network architecture called MCDFN, designed to analyze time series data comprehensively by incorporating multiple parallel channels within the network.
@@ 2 Literature Review
The application of DL in demand forecasting has seen significant advancements, especially in SCM. These innovations aim to enhance predictive accuracy and operational efficiency, which is crucial for maintaining robust and resilient SCs. This literature review explores recent studies from 2017 to 2025, highlighting the evolving methodologies and their implications in SCM. Recent research has introduced Temporal Fusion Transformers (TFT) as a powerful tool for handling multi-horizon forecasting problems (Lim et al.,, 2021). TFT integrates gating mechanisms, variable selection networks, and attention-based models to enhance accuracy and interpretability. This model has demonstrated superior performance across various applications by effectively managing static and dynamic inputs and providing insights into feature importance and relevant time steps. Terrada et al. discussed the integration of AI within SCM 4.0, emphasizing the role of DL methods such as ARIMA and LSTM (Terrada et al.,, 2022). Their findings show that DL methods significantly improve demand forecasting accuracy using historical transaction data, thus enhancing decision-making processes and balancing supply and demand more effectively.
Recent studies have explored hybrid models for demand forecasting (El-kenawy et al.,, 2024), advancements in machine learning for SC analytics (Abdollahzadeh et al.,, 2024), and XAI techniques in forecasting (El-Kenawy et al.,, 2024), while optimization frameworks for dynamic systems have also been investigated (El-Sayed et al.,, 2024). Attention-based models have shown great promise in demand forecasting by capturing long-term dependencies and relevant features from time series data. Google Research demonstrated that attention mechanisms could improve traditional DL models’ performance by focusing on the most critical parts of the input data, thereby increasing forecast accuracy and interoperability (Arık and Pfister,, 2021). Techniques such as LIME (Local Interpretable Model-agnostic Explanations) and SHAP (SHapley Additive exPlanations) have been applied to address the need for transparency in DL models. These methods help explain the predictions by showing the contribution of individual features, making DL models more user-friendly and trustworthy.
@@ 3.1.1 Dataset Characteristics
    Timestamp feature: This column records the exact date of each sale, providing a continuous chronological sequence vital for time series analysis. The timestamps are formatted in a standard date format (YYYY-MM-DD), facilitating easy parsing and manipulation for analytical purposes.
@@
    Quantity sold feature: This column reflects the units sold on each corresponding day. The sales quantities are recorded as integers, representing the daily demand for the product.
In addition to these primary features, external factors such as holidays were encoded as binary variables (e.g., $is\_holiday$) to capture periods of potentially higher sales. While the dataset does not include external factors like weather or macroeconomic indicators, including holiday information enhances the model’s ability to account for irregular demand spikes. Key observations are:
  • 1.
    Seasonality: Product demand exhibits seasonality, being lower at the beginning and end of each year and peaking in the middle.
  • 2.
    External Factors: Holidays significantly influence sales patterns, necessitating their inclusion as a binary feature.
  • 3.
    Missing Values: The dataset is complete, with no missing values reported. However, preprocessing steps were implemented to ensure robustness in case of future datasets with incomplete data.
@@
3.1.2 Preprocessing Steps
To prepare the dataset for training and evaluation, several preprocessing steps were undertaken:
@@ 3.1.2 Preprocessing Steps
Cyclic Encoding: To handle the periodic nature of time series data, DateTime features such as ‘Day of Month’, ‘Week’, ‘Month’, and ‘Year’ were extracted and transformed into cyclical features using sine and cosine transformations. This approach ensures that the periodicity is represented in a form that machine learning models can interpret accurately. Equations 14 illustrate the transformation process:
@@
Holiday Feature: Some products have higher sales during holidays. An ‘is_holiday’ feature was included to capture periods of potentially higher sales.
Normalization: Ensuring proper feature scaling is imperative in the preparatory stages of NN training. In sales forecasting, there is a scale difference between sales and other features. A prevalent method employed for this purpose is normalization, wherein the mean and standard deviation of each feature (except for the ‘is_holiday’ feature) are subtracted and divided, respectively. This procedure aids in standardizing the range of feature values, thereby enhancing the efficiency of network convergence. The ‘is_holiday’ feature is already in binary format, requiring no further normalization. Consequently, accurate training cannot be performed when inputs are used without being normalized. Moreover, sales follow Gaussian distribution inherently. After reviewing available normalization methods, the sale feature is standardized using the following equation.
@@
Handling Missing Values: Although the dataset used in this study is complete, preprocessing pipelines were designed to handle missing values in future datasets. Techniques such as forward-filling, backward-filling, or interpolation could be applied to address gaps in time-series data.
@@ 3.2 Deep Learning Models
In multi-step prediction tasks, the modeling paradigm shifts from forecasting individual future points to predicting a sequence of future values, requiring a more sophisticated approach. Unlike single-step forecasting, which predicts only a single future point, a multi-step model generates a coherent sequence of future observations. A prominent strategy for this is the “single-shot” model, where the entire sequence prediction is formulated and executed in a single inference step. In our research, we adopted the single-shot prediction methodology, enabling the model to forecast the complete time series trajectory simultaneously. This approach ensures that the demands across all output time steps are predicted concurrently, fostering a comprehensive understanding of the temporal dynamics in the dataset. Specifically, our multi-step models are trained using daily samples as input data to predict 30 days into the future based on a preceding window of 30 days’ historical data (Figures 6 and 7). We developed and benchmarked MCDFN against seven fine-tuned state-of-the-art DL models in SC demand forecasting, including BiLSTM, CNN, RNN, Vanilla LSTM, Stacked LSTM, FCN, and GRU.
@@ 3.2.1 BiLSTM
The activation functions, namely the hyperbolic tangent (tanh) and the sigmoid function for recurrent connections imbue the model with non-linearity and memory retention capabilities crucial for sequence modeling. A dropout layer with a dropout rate of 0.2 is employed after the LSTM layer to mitigate overfitting and enhance generalization performance. Following sequence processing, a flattening operation transforms the output into a one-dimensional vector, facilitating subsequent dense layer operations. A densely connected layer with an output dimensionality of 30 is introduced, initializing the kernel weights with zeros. This layer maps the extracted temporal features into the desired output space, effectively reconstructing the predicted sequence.
@@ 3.2.5 Stacked LSTM
The stacked LSTM model begins with an input layer for data shaped $(30,10)$, followed by two LSTM layers. Each LSTM layer is configured with 512 units, which was determined through hyperparameter tuning and employs ‘tanh’ activation and ‘sigmoid’ recurrent activation functions. Both LSTM layers use the ‘return_sequences=True’ parameter to pass the sequence output to the next layer. Between the LSTM layers, a Dropout layer is included to mitigate overfitting, with 0 dropout rate after tuning. After the LSTM layers, the model output is flattened and passed through a Dense layer, initialized with zeros, ensuring unbiased predictions. Finally, the output is reshaped to the desired dimensions (30, 1). Stacked LSTM was evaluated and achieved the best validation mean squared error of 28.17 after 58 trials, demonstrating its ability to capture complex temporal dependencies and provide accurate demand forecasts for an output sequence length of 30 steps.
3.2.6 FCN
Our Fully Connected Network (FCN) model starts with an input layer to handle data shaped $(30,10)$, followed by a dense layer with 512 units chosen through hyperparameter tuning. The activation function for this layer was selected as ‘tanh’ from the options ‘relu’ and ‘tanh’. A Dropout layer with a rate of 0.0 is included to prevent overfitting. Finally, the model outputs predictions through a dense layer with a linear activation function to match the dimensionality of the target feature. The FCN model was evaluated and achieved the best validation mean squared error of 27.79 after 54 trials.
The input layer of an FCN receives the input vector $\mathbf{x}$. For each hidden layer $l$ (where $l=1,2,\ldots,L$ with $L$ being the number of hidden layers), the output $\mathbf{h}^{(l)}$ is computed using equation 15.